[Hallc_running] [New Logentry] Hall C RC update: Friday, Jan. 14, 2022

gaskelld at jlab.org gaskelld at jlab.org
Fri Jan 14 08:30:01 EST 2022

Logentry Text:
There were more issues with an SHMS Q1 lead-flow controller yesterday during the day (left instead of right this time).  These were resolved near the end of day shift.  During day (when the hall was available) and swing shifts, Ops worked to fix BLM issues that were preventing reliable delivery of 42 uA to Hall C.  These issues were resolved by end of swing shift, and we have been running very efficiently during owl shift.

As of the end if run 14232 (ended at 6:20 am), we have accumulated 5859 good e-pi events (Q2=3.85, W=3.07, low epsilon,SHMS right setting) on LH2.  If we maintain the same running efficiency, we should be ready to switch to the aluminum dummy target by the end day/early swing shift.

There will be a brief RC meeting this morning at 9:30 am (note the earlier time):


Zoom link:



This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3971415
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220114/779176c1/attachment.html>

More information about the Hallc_running mailing list